OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 2x - 3y - z if x = -3, y = 2, and z = -2 10 -10 -14 14

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

no, it's -10 2(-3) - 3(2) - (-2) -6 - 6 + 2 -12 + 2 10

OpenStudy (anonymous):

-10* sorry

OpenStudy (anonymous):

I got a 100 on this test. it is -14

OpenStudy (anonymous):

- 10 not - 14

OpenStudy (anonymous):

O boy lol well ummmm

OpenStudy (anonymous):

i showed you the math, i don't know how bailey got 100 with the wrong answer. perhaps whoever graded it was wrong. i don't know, the answer is -10 because adding 2 to -12 will bring it down to -10 not -14

OpenStudy (anonymous):

this is the easiest thing on math u should be able to do it ur self this how u do it

OpenStudy (anonymous):

okay thank you everyone so much :)

OpenStudy (anonymous):

Reason.. I think this would be the answer.. 2x - 3y -z .. said if x = -3,y = 2 and z = -2 2 (-3) - 3(2) - (-2) - 6 - 6 + 2 >> -3x2 = -6 // 2x-3 = -6 // (-)(-) = + >> - ( -2) = 2 -10 Now the answer is -10..